1. For a business that makes advance provision for uncollectible Reuellt (a) Journalize the entries to record the following: (18 PTOS.) Record the adjusting entry at December 31, the end of the fiscal year, to provide for doubtful accounts. The accounts receivable account has a balance of $800,000, and the contra asset account (Allowance for Doubtful Account) before adjustment has a debit balance of $600. Analysis of the receivables indicates doubtful accounts of $18,000. In March of the following fiscal year, the $350 owed by Fronk Co. on account is written off as uncollectible Eight months later, $200 of the Fronk Co. account is reinstated and payment of that amount is received In October, $400 is received on the $600 owed by Dodger Co. and the remainder is written off as uncollectible. (1) (2) (3) (4) Based on the data in (a) (I) above, what is the net realizable value of the accounts receivable as reported on the balance sheet as of December 31? (3 PTOS.) (b)
Answer #1

(1) Dec,31                       Bad Debt Expense A/c (Dr)                       18600                                                      (18000+600)

                                         To Allowance for Doubtful Account(Cr)                18600

(2) March,Next year      Allowance for Doubtful Account(Dr)           350

                                      To Accounts Receivable A/c(Cr)                        350

(3)           (a)                 Accounts Receivable A/c (Dr)                      200

                                     To Allowance for Doubtful Account(Cr)                   200

               (b)                 Cash A/c (Dr)                                              200

                                     To Accounts Receivable A/c(Cr)                           200

(4) October                  Cash A/c(Dr)                                                400

                                    Allowance for Doubtful Account(Dr)             200

                                    To Accounts Receivable A/c(Cr)                             600

(B) Calculation of realizable value of AR

Allowance for doubtful debt balance= (600)+18600-350+200-200=17650

Accounts Receivable(Gross)=800000-350+200-200-600=799050

Net Realizable value of AR= AR(Gross) - Allowance for doubtful debt

                                           =799050-17650=781400